OpenStudy (anonymous):

solve for x: 1 over x minus 1 minus 1 over x plus 2

OpenStudy (anonymous):

\[\frac{ 1 }{ x }-1-\frac{ 1 }{ x }+2\]

OpenStudy (anonymous):

the minus one and plus 2 are with the x's

OpenStudy (anonymous):

Like \[\frac{ 1 }{ x-1 }-\frac{ 1 }{ x+2 }\]

OpenStudy (anonymous):

yes that's the problem

OpenStudy (anonymous):

can you do it